could you show me step by step how to figure this out Plz this is an example I don't want my homework done for me, I need to understand how to do it Thank you !
The problem is: Are the lines parallel,perpendicular or neither ? a) line 1: y=3x-5 line 2 : -3+12 or with something like line 1 : y=1/4x+1 and line 2 y=4x-5 any help would be great to get me to understand better what I am doing with this thank you

Answer by mananth(16946)

Are the lines parallel,perpendicular or neither ? a) line 1: y=3x-5 line 2 : -3+12 or with something like line 1 : y=1/4x+1 and line 2 y=4x-5

y=3x-5
y=3x+12
compare both equations with y= mx +b
in line 1 m=3, b=-5
line 2 m=3,b=12
slopes are same and the b is different so the lines are parallel.
If b is also same and m is also same then the equations represent the same line

y=1/4x+1
y=4x-5
here the slope is 1/4
and 4
The slopes are different
so lines are neither parallel, nor perpendicular They intersect
If the product of the slopes of the two lines = -1
then the lines are perpendicular